# When Speculation Spills Secrets: Side Channels via Speculative Decoding in LLMs

### v3

@@ Abstract
Deployed large language models (LLMs) often rely on speculative decoding, a technique that generates and verifies multiple candidate tokens in parallel, to improve throughput and latency. In this work, we reveal a new side-channel whereby input-dependent patterns of correct and incorrect speculations can be inferred by monitoring per-iteration token counts or packet sizes. We demonstrate that an adversary observing these patterns can fingerprint user queries with ¿90% accuracy across four speculative-decoding schemes, REST ($\sim$100%), LADE (up to 92%), BiLD (up to 95%), and EAGLE (up to 77.6%) and leak confidential datastore contents used for prediction at rates exceeding 25 tokens/sec. We evaluate the side-channel attacks in both research prototypes as well as the production-grade vLLM serving framework. To defend against these, we propose and evaluate a suite of mitigations, including packet padding and iteration-wise token aggregation.
@@ 1 Introduction
Large Language Models (LLMs) have transformed natural language processing (NLP), allowing machines to generate and understand human language at an unprecedented scale Vaswani et al. (2017); Devlin et al. (2019); Brown et al. (2020); Zhang et al. (2022); Le Scao et al. (2022); Touvron et al. (2023). LLMs typically generate text using auto-regressive decoding Touvron et al. (2023); OpenAI (2024), where the generation happens serially and each token depends on all the previous ones. Unfortunately, this serial process causes a significant bottleneck in the LLM response latency Miao et al. (2024); Fu et al. (2024) and under-utilizes the available hardware-level parallelism, limiting token generation throughput and latency.
@@
However, the adoption of speculation techniques is not without risks. Speculative execution in CPUs  Burton (1985); Hennessy & Patterson (2012), which inspired speculative decoding, has led to security vulnerabilities in processors, such as Spectre Kocher et al. (2019) and their variants, which exploit side-channels, i.e. timing variations due to mis-speculations that leak secret data accessed during mis-speculations. This raises the question, does speculative decoding in LLMs also introduce new risks to privacy? In this paper, we provide a study of the privacy risks of speculative decoding in LLMs, including leakage of private user inputs.
@@
Figure 1(a) shows a response by TinyLlama 1.1B Chat  Zhang et al. (2024a) with a speculative decoding technique, Lookahead Decoding (LADE) Fu et al. (2024), with correct (blue) and mis-speculated (red) tokens illustrated. On correct speculation, up to $n$ tokens are verified per iteration, as shown in Figure 1(b), while on mis-speculation, only one token is generated per iteration. In LLM applications streaming responses to the user, token by token, a network-based adversary, such as a malicious ISP or a compromised router Weiss et al. (2024), can observe network packet sizes corresponding to each iteration to deduce the number of tokens per iteration and learn the input-dependent pattern of correct and incorrect speculations. We target medical chatbots (e.g., Hippocratic AI Shah (2025)), where users interact by asking questions about symptoms or diseases, as this is a privacy-sensitive application where any information leakage (e.g., disease or symptoms of users) breaks privacy laws.
Exploits. We demonstrate privacy breaches based on speculation patterns, as shown in Figure 1(c). First, we show a query fingerprinting attack (e.g., leaking diseases or symptoms in medical chatbots) by a network adversary. By profiling speculation pattern fingerprints for prompt-response pairs offline, an adversary can then use compare the speculation pattern for an unknown query with the fingerprints to recover private user prompts. Across speculative decoding schemes (BiLD, REST, LADE) Fu et al. (2024); He et al. (2024); Kim et al. (2023), our prompt identification attacks reaches accuracies of $\sim$100% for REST, 92% for LADE, and 95% for BiLD with only 30 training samples. On a remote vLLM inference server Kwon et al. (2023), which supports speculation with EAGLE Li et al. (2024a), we similarly observe a high accuracy of 77.6%. We further show that when the actual prompts are unavailable, the adversary can train effective fingerprints using publicly available proxy datasets; e.g., by querying a medical chatbot with the top 50 most common diseases as stand-ins for real prompts. In this case, the attack still achieves 20–40% accuracy in recovering the disease or symptoms from user prompts, far exceeding random guessing (2%). Lastly, we show how adversaries crafting malicious inputs can also leak confidential data such as data-store contents used for predictions with REST He et al. (2024) at a rate of more than 25 tokens/second by observing correctly predicted tokens.
@@ 2.1 Prior Speculative Decoding Techniques
Speculation with Smaller Draft Models. Prior works propose using smaller, faster draft models to generate sequences of tokens speculatively Miao et al. (2024); Kim et al. (2023), and then verify a single sequence or a tree of such sequences in parallel using the larger target model. The draft model can be a smaller version of the model family or a pruned larger model Yan et al. (2025). In this paper, we demonstrate our attacks on BiLD Kim et al. (2023), as a example of this type of speculation.
Speculation via Self-Drafting. Recent works reuse the target model for producing predictions, eliminating the need for a separate draft model. Look-Ahead Decoding (LADE) Fu et al. (2024) caches previous output tokens and uses them to speculate future tokens via a key-value store populated with token sequence N-grams. These are populated using Jacobi decoding, by selecting random tokens from the input and generating successive chain of tokens. When this key token reappears in the output stream, LADE uses the stored N-grams for speculation. Medusa Cai et al. (2024) trains extra heads in the target model, each responsible for a token position in the draft sequence, for speculative tokens. EAGLE and EAGLE-2 Li et al. (2024a; b) perform auto-regressive decoding directly on the feature layers itself to generate speculative tokens, and are integrated in the vLLM inference serving platform. Our attacks target LADE (locally) and EAGLE (on a remote vLLM server) as representative examples, but our techniques are broadly applicable to other techniques as well.
Speculation Using Retrieval. REST He et al. (2024) generates predictions by retrieving data from a datastore containing relevant text or code, organized as a set of prefixes and continuations that can be potentially used as speculative tokens. If there are multiple candidates, REST organizes them in a tree, applying a tree attention mask during decoding with the target model.
Tree-Based Verification. SpecInfer Miao et al. (2024), SpecTr Sun et al. (2023) and REST He et al. (2024) generate multiple drafts of speculation per iteration, organized in a tree structure, to increase the length of successful verification for higher speedup. These drafts are organized with a custom tree attention mask to prevent inter-tree interactions and verified in parallel. In this work, we demonstrate our attacks on REST He et al. (2024), as a representative example of such techniques.
@@ 2.2 Side-Channel Attacks on LLMs
Debenedetti et al. (2024) pioneered side-channel attacks on LLMs. Their attack exploits variation in output token counts due to the tokenizer, to leak uncommon strings in the tokenizer vocabulary. Our work instead exploits variation in token count per iteration due to speculative decoding.
Weiss et al. (2024) identified a token-length side channel in streaming LLMs, where network attackers can infer token values by analyzing encrypted packet sizes. This attack reconstructs around 27% of the LLM outputs using the character length of each token. In our work, we similarly focus on a network-attacker observing the encrypted packet sizes, and leverage the variation in packet sizes based on speculative-decoding to identify user inputs. Moreover, we show mitigating the token-length side channel by padding tokens to the maximum token-size still results in packet size variations due to speculative decoding varying the output token counts per packet. We show that this can be used to fingerprint prompts and responses in our attacks with greater than 50% accuracy, even after the token-length side-channel is mitigated, in Appendix D.
Carlini & Nasr (2024) demonstrated timing attacks on LLM inference optimizations in commercial models like GPT-4 and Claude. Their attacks rely on timing variations due to these optimizations, i.e. variations in packet inter-arrival times, which is considerably susceptible to noise due to network congestion wand server loads. In comparison, our attacks exploit packet size variations due to inference optimizations, which are inherently more robust and independent of server or network load variations. We show the resilience of our attacks compared to this prior work in Section 4.7.
Song et al. (2024) and  Zheng et al. (2024) demonstrated timing side-channel attacks exploiting shared key-value (KV) and semantic caches to leak sensitive LLM inputs. Wiretapping LLMs Soleimani et al. (2025) also reveals private attributes in LLM-based medical and financial applications.
@@ 3 Threat Model
Like prior works Weiss et al. (2024), we focus on streaming LLMs, where the LLM server sends back responses to the user token-by-token. Here, we study the following attacker models.
Network Attacker. We assume a malicious ISP or a compromised router that can observe network packet sizes, similar to Weiss et al. (2024). The attacker’s goal is to leak private user queries. Although LLM response packets are encrypted, the packet size is still observable. The attacker can also observe the time between each packet to identify the iteration ID. We show how an adversary can fingerprint and leak private user queries based on this in Section 4.
Malicious User. Here we assume the attacker is a malicious user interacting with the LLM, similar to Debenedetti et al. (2024). The user can craft arbitrary inputs and inspect the output, and the tokens per iteration with the goal of leaking confidential data used by the LLM’s speculative decoding mechanism for generating predictions. We demonstrate this attack in Section 5.
@@ 4.2 Fingerprinting Using Speculation Patterns
To quantitatively validate this, we generated 2 traces each from 50 prompts taken from MedAlpaca dataset Han et al. (2023) (Human-LLM interactions asking questions on different diseases; see Appendix A) input to LADE, with temperature of 0.3, and computed the pairwise cosine similarity between them. Traces of the same prompt (or disease) have cosine similarity of 0.9–1, whereas traces from different prompts have cosine similarity between 0.4 - 0.8. This shows that the trace of tokens per iteration with speculative decoding can uniquely fingerprint and identify a prompt and response.
As packets are encrypted, the attacker cannot see the token counts directly. However, as shown in Figure 2(b), packet size strongly correlates with the number of tokens per iteration (Pearson Correlation Coefficient of 0.747 across all packets above). Thus, encrypted packet sizes serves as an effective proxy for the tokens per iteration in our fingerprinting attack. Note that exact token counts are not necessary; merely correlation of packet size with token counts suffices. Furthermore, any mitigation for prior token-size side-channels Weiss et al. (2024) by hiding individual token sizes is insufficient, as speculation itself introduces packet size variations due to varying token counts (see Appendix D).
@@ 4.3 Attack Design
Experiment 1. The attacker has exact knowledge of the complete set of queries, allowing the attacker to profile the exact queries offline and leak them out in the online phase. We choose a set of 50 prompts from a real-world human-LLM interactions dataset Han et al. (2023). These prompts ask a variety of questions about different diseases (the list of prompts is in Appendix A).
Experiment 2. Next, we ask the question, do the speculation patterns depend on structural patterns or deeper semantic meanings within the response? To analyze this better, we choose 50 prompts about diseases that all starts with “What are the symptoms of” from the same dataset, ensuring all prompts have a similar structure. With this, we seek to understand whether our attack can be used to leak deeper semantic meanings associated with the prompts (e.g., diseases of the prompts) even when all the prompts have similar structures.
Experiment 3. Assuming that speculation patterns rely on semantic meanings of responses, can the attacker use this to leak private information about the prompt (e.g., user’s disease) without the exact knowledge of the user’s query? While Experiment 1 and 2 assumed training and testing prompts to be the identical, here in Experiment 3, the attacker only has approximate knowledge of user queries during profiling. The attacker’s training set consists of queries that are semantically similar to user’s test-set queries (e.g., profiled query: “For how long should you take zolpidem?” and user’s query: “How long is zolpidem typically prescribed for?”). We use the prompts from Experiment 1 as our training prompts, but during testing, we rephrase the prompts using ChatGPT-4o by prompting it as ”Can you rephrase the following 50 questions while keeping the content in the question the same?”. We provide a more general experiment with an out-of-distribution training set in Section 4.8.
@@ 4.4 Experimental Setup
Target Models. We target four representative speculative decoding techniques: LADE Fu et al. (2024) and EAGLE Li et al. (2024a) using self-drafting, REST He et al. (2024) using retrieval and tree-verification, and BiLD Kim et al. (2023) using a smaller draft model. We run LADE with TinyLlama 1.1B Chat V1.0 Zhang et al. (2024a), REST with Vicuna 7B Chiang et al. (2023) as the target model and ShareGPT Aeala (2023) data-store of 120K conversations, and BiLD with Llama2 7B Touvron et al. (2023) and TinyLlama 1.1B Zhang et al. (2024a) as the target and draft model respectively. We run LADE and REST on NVIDIA RTX4090 (24GB) and BiLD on NVIDIA A100 (40GB) GPUs with Question-Answer tasks, to model a chatbot under attack.
Datasets and Classifier. For each experiment, we use 50 queries of varying complexity, length, and semantics (Appendix A lists the prompts). For the training set in the profiling phase of the attack, we run each query 5 to 30 times (with varying temperatures) and extract the packet size trace (5 to 30 traces per query) and train a Random Forest classifier. For the test set in the online phase, we use the corresponding 50 queries and 5 new traces per query, collecting 250 data points for evaluation, and report accuracy. For our attack, we use scikit-learn’s Random Forest with 150 estimators, max depth of 15, minimum 10 samples per split and 1 sample per leaf, and mean squared error as loss function.
Traces Per Query (TPQ)
5 10 20 30
Experiment 1
LADE 53.6% 63.2% 65.5% 68%
REST 91.2% 97.6% 100% 99.6%
BiLD 28.4% 41.6% 49.2% 52%
Experiment 2
LADE 33.2% 53.2% 64.8% 70.8%
REST 96.4% 97.6% 100% 99.6%
BiLD 58.8% 68% 75.2% 77.6%
Experiment 3
LADE 15.2% 16% 13.6% 18.8%
REST 35.2% 38.4% 40% 39.6%
BiLD 13.2% 20% 20.4% 20%
Table 1: Accuracy of the Query Fingerprinting Attack on LADE, REST, and BiLD with temperature of 0.8, using 5 to 30 Traces Per Query (TPQ) for training.
Temperature
0.3 0.6 0.8 1.0
Experiment 1
LADE 91.6% 85.6% 68% 61.2%
REST 100% 100% 99.6% 99.6%
BiLD 74.4% 59.2% 52% 34.8%
Experiment 2
LADE 90.8% 84.4% 70.8% 60.4%
REST 99.6% 100% 99.6% 100%
BiLD 95.2% 88.4% 77.6% 63.6%
Experiment 3
LADE 16.8% 17.6% 18.8% 19.6%
REST 35.2% 42% 39.6% 43.2%
BiLD 17.2% 15.6% 20% 16.4%
Table 2: Ablation Study. Attack accuracy of query fingerprinting as temperatures vary (0.3, 0.6, 0.8, 1.0), using 30 Traces Per Query (TPQ) for training.

### v4

@@ Abstract
Deployed large language models (LLMs) often rely on speculative decoding, a technique that generates and verifies multiple candidate tokens in parallel, to improve throughput and latency. In this work, we reveal a new side-channel whereby input-dependent patterns of correct and incorrect speculations can be inferred by monitoring per-iteration token counts or packet sizes. In evaluations using research prototypes and production-grade vLLM serving frameworks, we show that an adversary monitoring these patterns can fingerprint user queries (from a set of 50 prompts) with over 75% accuracy across four speculative-decoding schemes at temperature 0.3: REST (100%), LADE (91.6%), BiLD (95.2%), and EAGLE (77.6%). Even at temperature 1.0, accuracy remains far above the 2% random baseline—REST (99.6%), LADE (61.2%), BiLD (63.6%), and EAGLE (24%). We also show the capability of the attacker to leak confidential datastore contents used for prediction at rates exceeding 25 tokens/sec. To defend against these, we propose and evaluate a suite of mitigations, including packet padding and iteration-wise token aggregation.
@@ 1 Introduction
Large Language Models (LLMs) have transformed natural language processing (NLP), allowing machines to generate and understand human language at an unprecedented scale (Vaswani et al., 2017; Devlin et al., 2019; Brown et al., 2020; Zhang et al., 2022; Le Scao et al., 2022; Touvron et al., 2023). LLMs typically generate text using auto-regressive decoding (Touvron et al., 2023; OpenAI, 2024), where the generation happens serially and each token depends on all the previous ones. Unfortunately, this serial process causes a significant bottleneck in the LLM response latency (Miao et al., 2024; Fu et al., 2024) and under-utilizes the available hardware-level parallelism, limiting token generation throughput and latency.
@@
However, the adoption of speculation techniques is not without risks. Speculative execution in CPUs  (Burton, 1985; Hennessy & Patterson, 2012), which inspired speculative decoding, has led to security vulnerabilities in processors, such as Spectre (Kocher et al., 2019) and their variants, which exploit side-channels, i.e. timing variations due to mis-speculations that leak secret data accessed during mis-speculations. This raises the question, does speculative decoding in LLMs also introduce new risks to privacy? In this paper, we provide a study of the privacy risks of speculative decoding in LLMs, including leakage of private user inputs.
@@
Figure 1(a) shows a response by TinyLlama 1.1B Chat  (Zhang et al., 2024a) with a speculative decoding technique, Lookahead Decoding (LADE) (Fu et al., 2024), with correct (blue) and mis-speculated (red) tokens illustrated. On correct speculation, up to $n$ tokens are verified per iteration, as shown in Figure 1(b), while on mis-speculation, only one token is generated per iteration. In LLM applications streaming responses to the user, token by token, a network-based adversary, such as a malicious ISP or a compromised router (Weiss et al., 2024), can observe network packet sizes corresponding to each iteration to deduce the number of tokens per iteration and learn the input-dependent pattern of correct and incorrect speculations. We target medical chatbots (e.g., Hippocratic AI (Shah, 2025)), where users interact by asking questions about symptoms or diseases, as this is a privacy-sensitive application where any information leakage (e.g., disease or symptoms of users) breaks privacy laws.
Exploits. We demonstrate privacy breaches based on speculation patterns (Figure 1(c)). First, we show a query fingerprinting attack (e.g., leaking diseases or symptoms in medical chatbots) by a network adversary. By profiling speculation pattern fingerprints for prompt-response pairs offline, an adversary can then compare the speculation pattern for an unknown query with the fingerprints to recover private user prompts. Across speculative decoding schemes (BiLD, REST, LADE) (Fu et al., 2024; He et al., 2024; Kim et al., 2023), our prompt identification attacks (within a set of 50 queries) reaches accuracies of $\sim$100% for REST, 91.6% for LADE, and 95% for BiLD with a temperature of 0.3. On a remote vLLM inference server (Kwon et al., 2023), which supports speculation with EAGLE (Li et al., 2024a), we similarly observe a high accuracy of 77.6%. We further show that when the actual prompts are unavailable, the adversary can train effective fingerprints using publicly available proxy datasets; e.g., by querying a medical chatbot with the top 50 most common diseases as stand-ins for real prompts. In this case, the attack still achieves 20–40% accuracy in recovering the disease or symptoms from user prompts, far exceeding random guessing (2-6%). Lastly, we show how adversaries crafting malicious inputs can leak confidential data-store contents used for predictions with REST (He et al., 2024) at a rate of >>25 tokens/second by observing correctly predicted tokens.
@@ 2.1 Prior Speculative Decoding Techniques
Speculation with Smaller Draft Models. Prior works propose using smaller, faster draft models to generate sequences of tokens speculatively (Miao et al., 2024; Kim et al., 2023), and then verify a single sequence or a tree of such sequences in parallel using the larger target model. The draft model can be a smaller version of the model family or a pruned larger model (Yan et al., 2025). In this paper, we demonstrate our attacks on BiLD (Kim et al., 2023), as a example of this type of speculation.
Speculation via Self-Drafting. Recent works reuse the target model for producing predictions, eliminating the need for a separate draft model. Look-Ahead Decoding (LADE) (Fu et al., 2024) caches previous output tokens and uses them to speculate future tokens via a key-value store populated with token sequence N-grams. These are populated using Jacobi decoding, by selecting random tokens from the input and generating successive chain of tokens. When this key token reappears in the output stream, LADE uses the stored N-grams for speculation. Medusa (Cai et al., 2024) trains extra heads in the target model, each responsible for a token position in the draft sequence, for speculative tokens. EAGLE and EAGLE-2 (Li et al., 2024a; b) perform auto-regressive decoding directly on the feature layers itself to generate speculative tokens, and are integrated in the vLLM inference serving platform. Our attacks target LADE (locally) and EAGLE (on a remote vLLM server) as representative examples, but our techniques are broadly applicable to other techniques as well.
Speculation Using Retrieval. REST (He et al., 2024) generates predictions by retrieving data from a datastore containing relevant text or code, organized as a set of prefixes and continuations that can be potentially used as speculative tokens. If there are multiple candidates, REST organizes them in a tree, applying a tree attention mask during decoding with the target model.
Tree-Based Verification. SpecInfer (Miao et al., 2024), SpecTr (Sun et al., 2023) and REST (He et al., 2024) generate multiple drafts of speculation per iteration, organized in a tree structure, to increase the length of successful verification for higher speedup. These drafts are organized with a custom tree attention mask to prevent inter-tree interactions and verified in parallel. In this work, we demonstrate our attacks on REST (He et al., 2024), as a representative example of such techniques.
@@ 2.2 Side-Channel Attacks on LLMs
(Debenedetti et al., 2024) pioneered side-channel attacks on LLMs. Their attack exploits variation in output token counts due to the tokenizer, to leak uncommon strings in the tokenizer vocabulary. Our work instead exploits variation in token count per iteration due to speculative decoding.
(Weiss et al., 2024) identified a token-length side channel in streaming LLMs, where network attackers can infer token values by analyzing encrypted packet sizes. This attack reconstructs around 27% of the LLM outputs using the character length of each token. In our work, we similarly focus on a network-attacker observing the encrypted packet sizes, and leverage the variation in packet sizes based on speculative-decoding to identify user inputs. Moreover, we show mitigating the token-length side channel by padding tokens to the maximum token-size still results in packet size variations due to speculative decoding varying the output token counts per packet. We show that this can be used to fingerprint prompts and responses in our attacks with greater than 50% accuracy, even after the token-length side-channel is mitigated, in Appendix E.
(Carlini & Nasr, 2024) demonstrated timing attacks on LLM inference optimizations in commercial models like GPT-4 and Claude. Their attacks rely on timing variations due to these optimizations, i.e. variations in packet inter-arrival times, which is considerably susceptible to noise due to network congestion and server loads. In comparison, our attacks exploit packet size variations due to inference optimizations, which are inherently more robust and independent of server or network load variations. We show the resilience of our attacks compared to this prior work in Section 4.7.
(Song et al., 2024) and  (Zheng et al., 2024) demonstrated timing side-channel attacks exploiting shared key-value (KV) and semantic caches to leak sensitive LLM inputs. Wiretapping LLMs (Soleimani et al., 2025) also reveals private attributes in LLM-based medical and financial applications.
@@ 3 Threat Model
Like prior works (Weiss et al., 2024), we focus on streaming LLMs, where the LLM server sends back responses to the user token-by-token. Here, we study the following attacker models.
Network Attacker. We assume a malicious ISP or a compromised router that can observe network packet sizes, similar to  (Weiss et al., 2024). The attacker’s goal is to leak private user queries. Although LLM response packets are encrypted, the packet size is still observable. The attacker can also observe the time between each packet to identify the iteration ID. We show how an adversary can fingerprint and leak private user queries based on this in Section 4.
Malicious User. Here we assume the attacker is a malicious user interacting with the LLM, similar to  (Debenedetti et al., 2024). The user can craft arbitrary inputs and inspect the output, and the tokens per iteration with the goal of leaking confidential data used by the LLM’s speculative decoding mechanism for generating predictions. We demonstrate this attack in Section 5.
@@ 4.2 Fingerprinting Using Speculation Patterns
To quantitatively validate this, we generated 2 traces each from 50 prompts taken from MedAlpaca dataset (Han et al., 2023) (Human-LLM interactions asking questions on different diseases; see Appendix A) input to LADE, with temperature of 0.3, and computed the pairwise cosine similarity between them. Traces of the same prompt (or disease) have cosine similarity of 0.9–1, whereas traces from different prompts have cosine similarity between 0.4 - 0.8. This shows that the trace of tokens per iteration with speculative decoding can uniquely fingerprint and identify a prompt and response.
As packets are encrypted, the attacker cannot see the token counts directly. However, as shown in Figure 2(b), packet size strongly correlates with the number of tokens per iteration (Pearson Correlation Coefficient of 0.747 across all packets above). Thus, encrypted packet sizes serves as an effective proxy for the tokens per iteration in our fingerprinting attack. Note that exact token counts are not necessary; merely correlation of packet size with token counts suffices. Furthermore, any mitigation for prior token-size side-channels (Weiss et al., 2024) by hiding individual token sizes is insufficient, as speculation itself introduces packet size variations due to varying token counts (see Appendix E).
@@ 4.3 Attack Design
Experiment 1 - Exact Knowledge The attacker has exact knowledge of the complete set of queries, allowing the attacker to profile the exact queries offline and leak them out in the online phase. We choose a set of 50 prompts from a real-world human-LLM interactions dataset (Han et al., 2023). These prompts ask a variety of questions about different diseases (the list of prompts is in Appendix A).
Experiment 2 - Exact Knowledge (Constrained) Next, we ask the question, do the speculation patterns depend on structural patterns or deeper semantic meanings within the response? To analyze this better, we choose 50 prompts about diseases that all starts with “What are the symptoms of” from the same dataset, ensuring all prompts have a similar structure. With this, we seek to understand whether our attack can be used to leak deeper semantic meanings associated with the prompts (e.g., diseases of the prompts) even when all the prompts have similar structures.
Experiment 3 - Approximate Knowledge Assuming that speculation patterns rely on semantic meanings of responses, can the attacker use this to leak private information about the prompt (e.g., user’s disease) without the exact knowledge of the user’s query? While Experiment 1 and 2 assumed training and testing prompts to be the identical, here in Experiment 3, the attacker only has approximate knowledge of user queries during profiling. The attacker’s training set consists of queries that are semantically similar to user’s test-set queries (e.g., profiled query: “For how long should you take zolpidem?” and user’s query: “How long is zolpidem typically prescribed for?”). We use the prompts from Experiment 1 as our training prompts, but during testing, we rephrase the prompts using ChatGPT-4o by prompting it as ”Can you rephrase the following 50 questions while keeping the content in the question the same?”. We provide a more general experiment with an out-of-distribution training set in Section 4.8.
@@ 4.4 Experimental Setup
Target Models. We target four representative speculative decoding techniques: LADE (Fu et al., 2024) and EAGLE (Li et al., 2024a) using self-drafting, REST (He et al., 2024) using retrieval and tree-verification, and BiLD (Kim et al., 2023) using a smaller draft model. We run LADE with TinyLlama 1.1B Chat V1.0 (Zhang et al., 2024a), REST with Vicuna 7B (Chiang et al., 2023) as the target model and ShareGPT (Aeala, 2023) data-store of 120K conversations, and BiLD with Llama2 7B (Touvron et al., 2023) and TinyLlama 1.1B (Zhang et al., 2024a) as the target and draft model respectively. We run LADE and REST on NVIDIA RTX4090 (24GB) and BiLD on NVIDIA A100 (40GB) GPUs with Question-Answer tasks, to model a chatbot under attack.
Datasets and Classifier. For each experiment, we use 50 queries of varying complexity, length, and semantics (Appendix A lists the prompts). For our attack, we studied three types of classifiers - CNN, Gaussian Mixture Models (GMM) and Random Forest, and use Random Forest since it has the best accuracy. (See Figure 12 in Appendix G for the accuracy comparisons of Random Forest with CNN and GMM). For the training set in the profiling phase of the attack, we run each query 5 to 30 times (with varying temperatures) and extract the packet size trace (5 to 30 traces per query) to train our Random Forest classifier. For the test set in the online phase, we use the corresponding 50 queries and 5 new traces per query, collecting 250 data points for evaluation, and report accuracy. For our attack, we use scikit-learn’s Random Forest with 150 estimators, max depth of 15, minimum 10 samples per split and 1 sample per leaf, and mean squared error as loss function.
Figure 3: Accuracy of the query fingerprinting attack on LADE, REST, and BiLD with temperature of 0.8, using 5 to 30 Traces Per Query (TPQ) for training.